OpenStudy (anonymous):

The length of a rectangle is six times its width. If the area of the rectangle is 150in^2, find its perimeter.

OpenStudy (campbell_st):

draw a picture and label the sides, using the given information  next form the area equation \[x \times 6x = 150\] will 6x^2 = 150, x^2 = 25 only consider the positive value x = 5... width is 5 length is 30... this should be enough to find the perimeter